(1 + 1) REMPI SPECTROSCOPY OF NO COMPLEXES WITH N_{2}, CO AND CH_{4} VIA THE Ã STATE

\underline{JEROME LOZEILLE}, TIMOTHY G. WRIGHT, SOPHIA E. DAIRE and STUART D. GAMBLIN, Spectroscopy of Complexes and Radicals (SOCAR) Group, School of Chemistry, Physics and Environmental Science, University of Sussex, Falmer, Brighton, BN1 9QJ, UK.

The spectra of the Ã ← X̃ transition of the NO·X molecular complexes, where X = N_{2}, CO and CH_{4}, have been recorded using one-color (1 + 1) REMPI spectroscopy. The spectra exhibit vibronic features, which are interpreted as hindered internal rotation and the underlying end-over-end rotational structure. Some insights into the structure of the spectra are gleaned from ab initio calculations.